# Enhancing Brain Source Reconstruction through Physics-Informed 3D Neural Networks

### v1

@@ IV-B Data Generation
We employ a controlled synthetic EEG data generation process to facilitate model evaluation with known ground truth. To isolate the spatial reconstruction aspect of the problem, we simulate source activation at a single point in time. The source activations are modeled as 3-dimensional Gaussian bell curves centered at random locations $c$ within the brain. The width of each activation $\sigma$ is given by the standard deviation of the bell curve. This allows us to simulate both localized and diffuse patterns of brain activity. Sources are assigned random orientations $\vec{v}$ and their maximal amplitudes $a\sim\mathcal{N}(0,1)$ is sampled from a standard normal distribution. We sample $S\sim\mathcal{U}(1,N_{\mathrm{Activ}})$ distributed sources for a single data point. So given the center $c_{s}$, amplitude $a_{s}$ and direction $\vec{v}_{s}$, we can calculate the activation of each source $\textbf{x}_{i}$ by the sum of the $S$ sources:
@@
Here, $d$ measures the Euclidean distance between the 3D position of the source location $p_{i}$ and the center of active sources $c_{s}$. To generate the measurements $\textbf{y}=\textbf{L}\textbf{x}+\epsilon$, the brain activation is projected onto sensor space using the lead field matrices L described in the previous section. To introduce realistic challenges, zero-mean Gaussian noise $\epsilon\sim\mathcal{N}(0,\sigma_{\mathrm{Noise}}^{2}I)$ with varying signal-to-noise ratios (SNR), ranging from 0 dB to 30 dB, is applied to the measurements. More precisely, the noise standard deviation is given by:
$\sigma_{\mathrm{Noise}}=\frac{\sigma_{\mathrm{Signal}}}{\sqrt{10^{\mathrm{SNR}}}}.$ (5)
@@
  • SNR: Varied between 00 dB and $30$ dB to introduce realistic noise levels and challenge the model’s ability to reconstruct sources under different signal quality conditions.
This approach ensures a diverse dataset for training and evaluation. Fig. 2 illustrates the variety of generated source patterns.
@@ IV-C Evaluation Metrics
Figure 3: The ground truth activation and reconstruction from 3D-PIUNet , the fully connected Network and eLORETA for top a single active source with SNR of 00dB and bottom 3 active sources with SNR of $40$dB. 3D-PIUNet incorporates the learned data prior and refines the shape of an active region to match the ground truth. When multiple sources are active, it does focus on the strongest sources, missing weak activations in other parts of the brain.
@@ IV-C1 Normalized Mean Squared Error
Given the ground truth sources x and predicted sources $\hat{\textbf{x}}$, the Normalized Mean Squared Error (NMSE) measures the voxel- and direction-wise deviation normalized by the total activation of the ground truth:
$\text{NMSE}(\textbf{x},\hat{\textbf{x}})=\frac{\sum_{i=0}^{N}||\textbf{x}_{i}-\hat{\textbf{x}}_{i}||_{2}^{2}}{\sum_{i=0}^{N}||\textbf{x}_{i}||_{2}^{2}}$ (6)
By normalizing with the total activation, data samples with overall strong and weak brain activation are considered equally important along the validation dataset.
@@ IV-C2 Normalized Earth Mover’s Distance
To incorporate the spatial arrangement of brain sources and penalize larger misplacement of activations, Haufe et al. [61] proposed the use of the Earth Mover’s Distance (EMD) [62]. It measures the minimal transport needed to align the predicted and real source distributions, thereby capturing the location error. To unify different simulation settings, we propose to normalize the EMD by the EMD of the ground truth and an uniform distribution. To compute the EMD, we use the publicly available POT package [63].
@@ IV-C3 Weighted Cosine Distance
$\text{WCos}(\textbf{x},\hat{\textbf{x}})=1-\sum_{i}^{N}W_{i}\frac{\textbf{x}_{i}^{\top}\cdot\hat{\textbf{x}}_{i}}{||\textbf{x}_{i}||_{2}||\hat{\textbf{x}_{i}}||_{2}}$ (7)
The weight matrix $W_{i}=\frac{||\textbf{x}_{i}||_{2}}{\sum_{i}||\textbf{x}_{i}||_{2}}$ filters out non-active voxels and scales the importance proportional to the activation strength of the ground truth.
@@ IV-D Baselines
As a neural network competitor, we consider a fully connected network (FCN) with 4 fully connected layers and ReLU activation, upscaling the design of previous work [53]. The first layer maps from the $61$ measurements to $1024$ hidden dimensions, followed by two hidden layers with same dimensionality. The last layer maps from the $1024$ hidden dimensions to $4819\times 3=14457$ output channel. Note that this simple architecture already has around $17$ million parameters, more than the $11$ million parameters of 3D-PIUNet. We investigate different architecture sizes in Section V-F.
As classical baselines, we use eLORETA [10] as a minimum norm solution. Additionally, we compare against Lasso [20] with a group lasso penalty across three spatial dimensions, although we expect sparse methods to perform poorly with wide sources. Hyperparameters for these methods are optimized on the training data via a grid search.
@@ V-A Visual Results
In Fig. 3, we visually compare source predictions for eLORETA and 3D-PIUNet for a single active source with 20mm extent reconstructed from single timestep EEG measurements for a single active source with 00dB noise (top) and three active sources with $40$db noise (bottom). 3D-PIUNet reconstruction demonstrates a clear refinement of the eLORETA solution, more accurately capturing the spatial extent and peak activation of the ground truth.
@@ V-B Effect of Signal to Noise Ratio
Figure 4: Evaluation on the validation set across different Signal-to-Noise Ratios (SNR). The shaded area represents the standard deviations calculated from 5 trained models, emphasizing the consistency and reliability of the observed trends. Left: Normalized EMD decreases as SNR increases, with 3D-PIUNet consistently showing the lowest values, indicating superior performance in minimizing distribution discrepancies. Center: For low SNR, both neural network approaches substantially outperform eLORETA. The discrepancy shrinks as SNR increases, with 3D-PIUNet consistently achieving lower mean squared error. Right: Weighted Cosine Distance remains high across all SNR levels for the neural networks, suggesting robust similarity in orientation, while eLORETA shows a significant improvement with increasing SNR.
In this experiment, we systematically vary the SNR from high noise setting with 00dB to nearly noise-free observations of $40$dB. The results are depicted in Fig. 4. We evaluated reconstruction performance using proposed metrics from section 3: Earth Mover’s Distance (EMD) to quantify differences in spatial distributions, Mean Squared Error (MSE) for overall amplitude accuracy, and weighted cosine similarity to assess orientation similarity. Both neural network-based approaches (FCN and 3D-PIUNet) consistently outperform eLORETA across all metrics and SNR levels. Notably, 3D-PIUNet exhibited superior performance compared to the FCN, particularly at SNR levels above 10 dB. This suggests that the U-Net architecture, with its ability to learn complex spatial patterns and non-linear relationships, is better equipped to leverage the additional information available in higher SNR data.
@@ V-C Size of Sources
Figure 5: 3D-PIUNet outperforms all baselines for different spatial extent of a single source with $5$dB noise.
In Fig. 5, we show model performance on different-sized sources for a single active source with an SNR of $20$dB. 3D-PIUNet improves on the solution of eLORETA and outperforms the Fully Connected Network with respect to all metrics. For small active sources with an extent of 10mm, we observe a low EMD, indicating that 3D-PIUNet predicts activation close to the correct source, even outperforming the sparse method Lasso. However, the high Normalized MSE suggests that none of the approaches perfectly matches the source on a voxel level. This trade-off may arise from the model’s focus on accurately localizing sources, even when their exact amplitude is harder to capture at this resolution. Notably, 3D-PIUNet demonstrates robustness to variations in source extent, effectively handling both sparse and spread sources despite being trained on a general dataset. This suggests the model has learned a versatile representation of source activity patterns.
@@ V-D Number of Sources
Figure 6: We show the performance of increasing number of sources for $20$dB Noise with varying extension. Even though the performance drops with more than 3 active regions, 3D-PIUNet remains superior to the other methods for up to 5 active sources.
In Fig. 6, we analyze model performance for different numbers of active sources for the medium noise setting of 20dB and sources with random extent between $10$ and $80$mm. 3D-PIUNet consistently outperforms the FCN across all metrics, but shows a decrease in performance as the number of sources increases, particularly in terms of the normalized EMD. Interestingly, the performance of eLORETA remains relatively constant regardless of the number of active sources. This is likely because eLORETA is a linear method that distributes the estimated source activity smoothly across the entire source space. As the number of sources increases, eLORETA continues to spread the activity, leading to a consistent but less accurate reconstruction compared to the neural network models, which can learn more complex, non-linear relationships between sources and sensor data. As we consider only a single time step the signal of the sources overlap and the demixing becomes a very hard task, even when a data prior is learned. In future work, we aim to add the time dimension with varying source amplitudes, providing additional information necessary for separating the different regions.
@@ V-E Transferability to Real Data
Figure 7: This figure illustrates the predicted source activation for 20 ensembled stimuli from the THINGS-EEG dataset. Both panels display activation peaks of approximately 30 ms and around 110 ms post-stimulus exposure. In panel (a), the eLORETA solution shows a diffuse activation pattern, indicating a broader but less specific response within the visual cortex. In contrast, panel (b) demonstrates that 3D-PIUNet predicts localized and concentrated activation, specifically highlighting the visual cortex region. Additionally, the temporal response in eLORETA appears smoother and less defined, whereas 3D-PIUNet provides sharper, more distinct activation peaks, suggesting its superior ability to pinpoint the precise timing of neural activations.
@@
Using the pretrained model from last sections, we validate 3D-PIUNet on real data, utilizing the THINGS-EEG2 dataset [64]. This large-scale visual dataset followed the Rapid Serial Visual Presentation (RSVP) paradigm [65, 66] where images are displayed for 100 ms each, followed by a 100 ms blank screen in a series of 20 images. As the full dataset is large, we focus on the test set of subject 10, containing 20 trials of EEG recording for each image stimulus. We applied multivariate noise normalization [67], following the original work.
@@
The THINGS-EEG2 dataset does not include MRI data for the subjects. Consequently, we use the same head geometry from section IV-A and recomputing the forward and pseudo-inverse for the new sensor layout.To enhance the signal-to-noise ratio, we apply ensemble averaging [68], combining EEG signals from $20$ instances where the same image was presented to the subject. We apply eLORETA and 3D-PIUNet, to the preprocessed EEG data to estimate the source activations. As both are only spatial models, we predict for each time step independently. Note that 3D-PIUNet is trained only on the artificial data and applied to the different sensor layout without any fine-tuning by replacing just the pseudo inverse.
@@
The predicted source activation of a single image condition is shown in Fig. 7. We observe two main activations in the visual cortex, the first around 30 ms and the second 130 ms after stimulus exposure. As the experiments follow the RSVP paradigm, we assume, that the first peak corresponds to the previous image while the peak around 130 ms is consistent with the image correlation results from the original paper [64].
@@
The spatial and temporal characteristics of the activations differ significantly between the two methods. For eLORETA, the spatial activation pattern appears diffuse, while 3D-PIUNet is more localized, precisely highlighting the visual cortex. This concentrated activation suggests that 3D-PIUNet is better at pinpointing specific regions of neural activity. Additionally, the temporal response of eLORETA is smoother and less defined. While this can be beneficial for capturing sustained neural activity, it might miss rapid changes or peaks in activation. Despite being trained only on a single time step synthetic data, 3D-PIUNet provides sharper, more distinct activation peaks, indicating a superior ability to detect precise timing of neural responses, which is crucial for understanding fast neural dynamics. It also matches the expected visually evoked potential [69], showing three distinct peaks around 100 ms after stimulus. This verifies the the effectiveness of the proposed method also for real-world data.
@@ V-F Computational Efficiency
TABLE I: Comparison of different model configurations with respect to runtime and EMD for 200 single source activations with $20$ dB SNR and varying source extent. While 3D-PIUNet is slower than the Fully Connected Network, its runtime is still an order of magnitude faster than the iterative method Lasso.
Model Depth Width Parameter Runtime EMD
eLORETA 3 1005 0.0 Mio 0.010s 0.60
Lasso 0 0 0.1 Mio 21.645s 0.70
FCN 3 512 8.0 Mio 0.006s 0.50
FCN 3 1024 17.0 Mio 0.004s 0.44
FCN 4 1024 18.0 Mio 0.004s 0.46
FCN 3 4096 93.0 Mio 0.003s 0.33
FCN 3 16384 774.0 Mio 0.003s 0.30
3D-PIUNet 3 8 0.7 Mio 0.191s 0.17
3D-PIUNet 3 32 11.1 Mio 0.415s 0.17
3D-PIUNet 4 32 19.1 Mio 0.866s 0.16
3D-PIUNet 3 64 44.2 Mio 2.192s 0.16
While 3D-PIUNet offers superior reconstruction accuracy, it is important to consider its computational demands. Training 3D-PIUNet requires approximately 5 hours on a single A100 GPU, compared to 2 hours for the FCN. However, the primary concern for practical applications is inference time. In our evaluation on a validation set of 200 samples (Table I), both eLORETA and the Fully Connected Network perform inference for 200 samples in a fraction of seconds. 3D-PIUNet, while slower, still completed inference in less than a second, notably faster than Lasso. While scaling up the FCN improves its performance, there remains a big gap even to the smallest version of 3D-PIUNet, demonstrating the parameter efficiency of the convolutional network.

### v2

@@ IV-B Data Generation
Similar to previous work [35], we employ a controlled synthetic EEG data generation process to facilitate model evaluation with known ground truth. To isolate the spatial reconstruction aspect of the problem, we simulate source activation at a single point in time. The source activations are modeled as 3-dimensional Gaussian bell curves centered at random locations $c$ within the brain[32]. The width of each activation $\sigma$ is given by the standard deviation of the bell curve. This allows us to simulate both localized and diffuse patterns of brain activity. Sources are assigned random orientations $\vec{v}$ and their maximum amplitudes $a\sim\mathcal{N}(0,1)$ are sampled from a standard normal distribution. For each training example, we randomly select a number of simultaneously active distributed sources $S$, sampled from a uniform distribution $\mathcal{U}(1,N_{\mathrm{Activ}})$, where $N_{\mathrm{Activ}}$ defines the maximum number of active sources allowed. For each of these sources, we sample the center $c_{s}$, amplitude $a_{s}$ and unit norm direction $\vec{v}_{s}$. We calculate the activation of the source location $\textbf{x}_{i}$ by the sum of the $S$ active sources:
@@
Here, $d$ measures the Euclidean distance between the 3D position $p_{i}$ of the source location $x_{i}$ and the center of each active source $c_{s}$. To generate the measurements $\textbf{y}=\textbf{L}\textbf{x}+\epsilon$, the brain activation is projected onto sensor space using the lead field matrices L described in the previous section. To introduce realistic challenges, zero-mean Gaussian noise $\epsilon\sim\mathcal{N}(0,\sigma_{\mathrm{Noise}}^{2}I)$ with varying signal-to-noise ratios (SNR), ranging from 0 dB to 30 dB, is applied to the measurements. More precisely, the noise standard deviation is given by:
$\sigma_{\mathrm{Noise}}=\frac{\sigma_{\mathrm{Signal}}}{\sqrt{10^{0.1\mathrm{SNR}}}}.$ (6)
@@
  • SNR: Varied between 0 dB and $30$ dB to introduce realistic noise levels and challenge the model’s ability to reconstruct sources under different signal quality conditions.
Fig. 3 illustrates the variety of generated source patterns. While this controlled activation model allows for a systematic evaluation of spatial source reconstruction, the addition of temporal dynamics or more realistic brain activation, e.g. including brain parcellation maps[13, 17], could further improve the learned priors.
@@ IV-C Evaluation Metrics
Figure 4: The ground truth activation and reconstruction from 3D-PIUNet , the fully connected Network and eLORETA for top a single active source with SNR of 0dB and bottom 3 active sources with SNR of 4040dB. 3D-PIUNet incorporates the learned data prior and refines the shape of an active region to match the ground truth. When multiple sources are active, it does focus on the strongest sources, missing weak activations in other parts of the brain.
@@ IV-C1 Normalized Mean Squared Error
Given the ground truth sources $\textbf{x}\in\mathbb{R}^{3N}$ and predicted sources $\hat{\textbf{x}}\in\mathbb{R}^{3N}$ and $\textbf{x}_{i}\in\mathbb{R}^{3}$ defined as the oriented activation of source $i$, the Normalized Mean Squared Error (NMSE) measures the voxel- and direction-wise deviation normalized by the total activation of the ground truth:
[…]
By normalizing with the total activation per sample, we ensure that all samples contribute equally to the validation loss, regardless of their overall activation strength, while preserving relative source amplitudes within each sample.
@@ IV-C2 Normalized Earth Mover’s Distance
To incorporate the spatial arrangement of brain sources and penalize larger misplacement of activations for distributed dipoles, Haufe et al. [31] proposed the use of the Earth Mover’s Distance (EMD) [61]. It measures the minimal transport, based on the Euclidean distance in 3D space, needed to align the predicted and real source activations, thereby capturing the location error. Since brain sources often consist of multiple, spatially extended activation areas, we compute EMD between the predicted $\hat{\textbf{x}}$ and ground truth x source activations after normalizing them into probability distributions:
[…]
The EMD is then computed as the solution to the following constrained optimization problem:
[…]
$\text{s.t. \qquad}F$ $\geq 0;\quad\sum_{j}^{N}F_{ij}=\textbf{x}^{A}_{i};\quad\sum_{i}^{N}F_{ij}=\hat{\textbf{x}}^{A}_{j}$ (10)
[…]
$\text{NEMD}(\textbf{x},\hat{\textbf{x}})=\frac{\text{EMD}(\textbf{x},\hat{\textbf{x}})}{\text{EMD}(\textbf{x},\textbf{u})}$ (11)
where $\textbf{u}=\frac{1}{N}\mathbf{1}$ represents a uniform activation distributed equally across all potential source locations. This normalization step accounts for variations in source activation area due to multiple sources or large source extents, preventing automatic reductions in EMD caused by distributed activations rather than true mislocalization.
@@ IV-C3 Weighted Cosine Distance
[…]
The relative amplitude $\textbf{x}^{A}_{i}$ filters out non-active voxels and scales the importance proportional to the activation strength of the ground truth.
@@ IV-D Baselines
As a neural network competitor, we restrict our comparison to models that operate on a single time step, as our method does not leverage temporal context. Therefore, we consider the ConvDip model from [35] and a fully connected network (FCN) with 4 fully connected layers and ReLU activation, upscaling the design of previous work [36]. The first layer maps from the $61$ measurements to $C=1024$ hidden dimensions, followed by two hidden layers with the same dimensionality. The last layer maps from the $1024$ hidden dimensions to $4819\times 3=14457$ output channel. Note that this simple architecture already has around $17$ million parameters, more than the $11$ million parameters of 3D-PIUNet. We investigate different architecture sizes in Section V-F. As classical baselines, we use eLORETA [55] as a minimum norm solution. Additionally, we compare against Lasso [72] with a group lasso penalty across three spatial dimensions, although we expect sparse methods to perform poorly with wide sources. Hyperparameters for these methods are optimized on the training data via a grid search. We evaluate all approaches on the new head geometry, with a distinct number of possible source locations. Since ConvDip and FCN have a fixed output dimensionality, we morph their predictions to the new geometry using Symmetric Diffeomorphic Registration (SDR) [1]. In contrast, 3D-PIUNet, similar to the classical approaches eLORETA and Lasso, does not require SDR and instead adapts naturally by replacing the pseudo inverse. To ensure a realistic evaluation and avoid the inverse crime, this pseudo-inverse is computed using the conductivity values employed during training, rather than those used for data generation (see Sec. IV-A).
@@ V-A Visual Results
In Fig. 4, we visually compare source predictions for eLORETA and 3D-PIUNet for a single active source with 20mm extent reconstructed from single timestep EEG measurements for a single active source with 0dB noise (top) and three active sources with 4040db noise (bottom). 3D-PIUNet reconstruction demonstrates a clear refinement of the eLORETA solution, more accurately capturing the spatial extent and peak activation of the ground truth.
@@ V-B Effect of Signal to Noise Ratio
Figure 5: Evaluation on the validation set across different Signal-to-Noise Ratios (SNR). The shaded area represents the standard deviations calculated from 5 trained models, emphasizing the consistency and reliability of the observed trends. Left: Normalized EMD decreases as SNR increases, with 3D-PIUNet consistently showing the lowest values, indicating superior performance in minimizing distribution discrepancies. Center: For low SNR, the neural network approaches substantially outperform eLORETA. The discrepancy shrinks as SNR increases, with 3D-PIUNet consistently achieving lower mean squared error. Right: Weighted Cosine Distance remains low across all SNR levels for the neural networks, suggesting robust similarity in orientation, while eLORETA shows a significant improvement with increasing SNR.
In this experiment, we systematically vary the SNR from high noise setting with 0dB to nearly noise-free observations of 4040dB, with the latter containing less noise than the maximum SNR of 3030dB seen during training. The results are depicted in Fig. 5. We evaluated reconstruction performance using proposed metrics from section 4: Earth Mover’s Distance (EMD) to quantify differences in spatial distributions, Mean Squared Error (MSE) for overall amplitude accuracy, and weighted cosine similarity to assess orientation similarity. The neural network-based approaches (FCN, ConvDip and 3D-PIUNet) consistently outperform eLORETA and Lasso across all metrics and SNR levels with an exception of ConvDip for low noise scenarios. Since classical approaches reconstruct the measurements directly, they struggle with low SNR, while their performance converges toward that of neural networks as noise diminishes. Notably, 3D-PIUNet exhibited superior performance compared to the FCN, particularly at SNR levels around 10 dB. This suggests that the U-Net architecture, with its ability to learn complex spatial patterns and non-linear relationships, is better equipped to leverage the additional information available in higher SNR data. The slight increase in variance observed at very high SNRs is likely due to the mismatch between the forward model used for data generation and the pseudo-inverse used for initialization. Nonetheless, the overall performance of 3D-PIUNet remains consistently strong across all noise levels.
@@ V-C Size of Sources
Figure 6: 3D-PIUNet outperforms all baselines for different spatial extent of a single source with 55dB noise.
In Fig. 6, we show model performance on different-sized sources for a single active source with an SNR of 55dB. 3D-PIUNet improves on the solution of eLORETA and outperforms the FCN with respect to all metrics. For small active sources with an extent of 10mm, we observe a low EMD, indicating that 3D-PIUNet predicts activation close to the correct source, even outperforming the sparse method Lasso, which suffers from the small SNR as shown in the previous section. However, the high Normalized MSE suggests that none of the approaches perfectly matches the source on a voxel level. This trade-off may arise from the model’s focus on accurately localizing sources, even when their exact amplitude is harder to capture at this resolution. Notably, 3D-PIUNet demonstrates robustness to variations in source extent, effectively handling both sparse and spread sources despite being trained on a general dataset. This suggests the model has learned a versatile representation of source activity patterns.
@@ V-D Number of Sources
Figure 7: We show the performance of increasing number of sources for 2020dB Noise with varying extension. Even though the performance drops with more than 3 active regions, 3D-PIUNet remains superior to the other methods for up to 5 active sources.
In Fig. 7, we analyze model performance for different numbers of active sources for the medium noise setting of 20dB and sources with random extent between $10$ and 8080mm. 3D-PIUNet consistently outperforms the FCN across all metrics, but shows a decrease in performance as the number of sources increases, particularly in terms of the normalized EMD. Interestingly, the performance of eLORETA remains relatively constant regardless of the number of active sources. This is likely because eLORETA is a linear method that distributes the estimated source activity smoothly across the entire source space. As the number of sources increases, eLORETA continues to spread the activity, leading to a consistent but less accurate reconstruction compared to the neural network models, which can learn more complex, non-linear relationships between sources and sensor data. As we consider only a single time step the signal of the sources overlap and the demixing becomes a very hard task, even when a data prior is learned. In future work, we aim to add the time dimension with varying source amplitudes, providing additional information necessary for separating the different regions.
@@ V-E Transferability to Real Data
Figure 8: This figure illustrates the average predicted source activation for 20 ensembled stimuli from the THINGS-EEG dataset. Both panels display activation peaks of approximately 30 ms and around 110 ms post-stimulus exposure. In panel (a), the eLORETA solution shows a diffuse activation pattern, indicating a broader but less specific response within the visual cortex. In contrast, panel (b) demonstrates that 3D-PIUNet predicts localized and concentrated activation, specifically highlighting the visual cortex region.
@@
Using the pretrained model from the last sections, we validate 3D-PIUNet on real data, utilizing the THINGS-EEG2 dataset [23]. This large-scale visual dataset followed the Rapid Serial Visual Presentation (RSVP) paradigm [39, 26] where images are displayed for 100 ms each, followed by a 100 ms blank screen in a series of 20 images. As the full dataset is large, we focus on the test set of subject 10, containing 20 trials of EEG recording for each image stimulus. We applied multivariate noise normalization [27], following the original work.
@@
The THINGS-EEG2 dataset does not include MRI data for the subjects. Consequently, we use the same head geometry from section IV-A and recomputing the forward and pseudo-inverse for the new sensor layout. To enhance the signal-to-noise ratio, we apply ensemble averaging [66], combining EEG signals from $20$ instances where the same image was presented to the subject. We apply eLORETA and 3D-PIUNet, to the preprocessed EEG data to estimate the source activations. As both are only spatial models, we predict each time step independently. Note that 3D-PIUNet is trained only on the artificial data and applied to the different sensor layout without any fine-tuning by replacing just the pseudo inverse. Classical end-to-end approaches would require at least retraining of the input layer to account for the changed input dimensionality.
@@
The predicted source activation averaged over the 200 test image conditions is shown in Fig. 8. We observe two main activations in the visual cortex, the first around 30 ms and the second 110 ms after stimulus exposure. The first sharper peak after 25ms appears unusually early. As the experiments follow the RSVP paradigm, we assume that this first peak corresponds to the blank screen, which triggers a more uniform brain response. The second peak aligns well with canonical visual evoked responses, typically occurring between 75-140ms [10], and is consistent with the image-related cortical activity reported by Gifford et al. [23].
@@
The spatial characteristics of the activations differ significantly between the two methods. For eLORETA, the spatial activation pattern appears diffuse, while 3D-PIUNet is more localized, precisely highlighting the visual cortex. This concentrated activation suggests that 3D-PIUNet is better at pinpointing specific regions of neural activity. The temporal activation is less pronounced but displays a similar pattern to eLORETA. This verifies the effectiveness of the proposed method also for real-world data.
@@ V-F Computational Efficiency
TABLE I: Comparison of different model configurations with respect to runtime and EMD for 2000 single source activations with $20$ dB SNR and varying source extent. While 3D-PIUNet is slower than the Fully Connected Network, its runtime is still sufficiently fast for real time analysis. The number of parameters is given in million.
Model Depth Channels $C$ Parameter Runtime EMD
eLORETA - - 0.0M 0.003s 0.58
Lasso - - 0.0M 8.178s 0.72
FCN 3 512 8.0M 0.006s 0.52
FCN 3 1024 17.0M 0.006s 0.47
FCN 4 1024 18.0M 0.007s 0.49
FCN 3 4096 93.0M 0.006s 0.40
FCN 3 16384 774.0M 0.007s 0.38
ConvDip 3 512 8.0M 0.633s 0.60
3D-PIUNet 3 8 0.7M 0.967s 0.21
3D-PIUNet 3 32 11.1M 3.234s 0.20
3D-PIUNet 4 32 19.1M 3.230s 0.19
3D-PIUNet 3 64 44.2M 6.721s 0.20
While 3D-PIUNet offers superior reconstruction accuracy, it is important to consider its computational demands. Training 3D-PIUNet requires approximately 5 hours on a single A100 GPU, compared to 2 hours for the FCN. However, the primary concern for practical applications is inference time. In our evaluation on a validation set of 2000 samples (Table I), both eLORETA and the Fully Connected Network perform inference for 2000 samples in a fraction of seconds. 3D-PIUNet, while slower, still completed inference a few seconds, notably faster than Lasso. To further examine model efficiency, we compared different hyperparameter settings. While scaling up the FCN improves its performance, there remains a big gap even to the smallest version of 3D-PIUNet, demonstrating the parameter efficiency of the convolutional network. For 3D-PIUNet, we varied not only the number of channels $C$, but also the depth, defined as the number of encoder-decoder resolution levels (i.e., downscaling and upscaling stages). Overall, 3D-PIUNet shows stable performance across different hyperparameter settings.
V-G Loss function Overview
Figure 9: Comparison of training 3D-PIUNet with different loss functions. The proposed $\ell_{1}$ loss outperforms the $\ell_{2}$ loss for a single active source. Adding a data reconstruction loss does harm the performance.
In this section, we evaluate the proposed 3D-PIUNet trained with both $\ell_{2}$ and a data reconstruction loss on single-source prediction across three different noise levels, as shown in Fig. 9. The reconstruction loss is based on the forward model and transfers the prediction back to measurement space:
[…]
However, due to the noisy measurements, this reconstruction loss drives the network’s predictions away from the ground truth, degrading performance. Overall, our results indicate that the $\ell_{1}$ loss in source space is the most effective for training the neural network.